OpenStudy (anonymous):

 Teresa says the two expressions below are equivalent:     2(5x-3)+6x and 16x-3    Is Teresa correct? Explain why or why not.  

OpenStudy (anonymous):

2(5x-3) = 2*5x + (2*(-3)) = 10x-6 and we had 6x more so 10x+6x-6 is going to be 16x-6 16x-6 is not equal to 16x-3 To give an explanation : If we assume that they are equal 16x-6 =16x-3 16x's gonna reset theirselves , so we will have the equation of -6=-3 which is not TRUE.

OpenStudy (anonymous):

thank you :))